OpenStudy (anonymous):

Find the unknown side length for the triangle. Area = 18 cm² A. 9 cm B. 14 cm C. 18 cm D. 36 cm posting pic below

OpenStudy (anonymous):



OpenStudy (anonymous):

@jango_IN_DTOWN

OpenStudy (anonymous):

The equation for the area of a triangle: A =1/2 x b x h

OpenStudy (anonymous):

From that, you can work out the bottom length. Try to do that first

OpenStudy (jango_in_dtown):

Its a right angled triangle?

OpenStudy (anonymous):

yeah the pic is above @jango_IN_DTOWN

OpenStudy (jango_in_dtown):

ok

OpenStudy (jango_in_dtown):

let the base be b then by the given data 18=4b/2 implies 2b=18 implies b=9

OpenStudy (anonymous):

what is the answer im trying to pass test @jango_IN_DTOWN

OpenStudy (jango_in_dtown):

now the hypotenuse is given by \[\sqrt{b^2+h^2}=\sqrt{9^2+4^2}=\sqrt{97}\]

OpenStudy (jango_in_dtown):

the options are not matching

OpenStudy (rccb):

just take 4, multiply it by 9, then divide it by 2, if that doesn't work try B just take 4, multiply it by 14, then divide it by 2, and continue until you get 18